OpenStudy (anonymous):

solve the following systems 6x + 3y = 6 8x + 5y = 12 please show work if you can!

OpenStudy (mathmale):

Let's use the 'elimination by addition and subtraction' method. Take the first equation 6x + 3y = 6 and divide every one of these three terms by -3. what do you get? Next, plese take the second equation: 8x + 5y = 12, and divide all of these three terms by 5. What do you get? type out your results.

OpenStudy (triciaal):

to use substitution divide equation 1 by 3 to get 2x + y = 2 y = 2-2x substitute in equation 2 8x + 5(2-2x) = 12 solve for x use this value of x in y = 2-2x good practice to check your solution by putting in either of original equations

OpenStudy (triciaal):

@mathmale to use that method it is always better to "match" to coefficient of the variable you want to eliminate. you do not have to let the coefficient = 1 and have to work with fractions etc. for this problem multiply the first equation by 5 so when you subtract the y's cancel out 10x -8x = 10 - 12

OpenStudy (anonymous):

To continue Triciaal's excellent answer, you could have, instead, 1. multiply the entire first equation by 5 (the coefficient of y in the second equation) 2. multiply the entire second equation by 3 (the coefficient of y in the first equation) This gives us: 30x + 15y = 30 24x + 15y = 36 Subtract, and we get 6x = -6 and I bet you can finish it from here. The nice thing about math is that there are usually multiple ways to get to the same answer.